OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 2197 ft3?

OpenStudy (welshfella):

x^3 = 2197 use ur calculator to find the cube root of 2197

OpenStudy (welshfella):

if you have.nt got a cube root key use 2197 ^ (0.3333)

OpenStudy (anonymous):

i dont get it i am in 6th grade remember i dont what a root key

OpenStudy (welshfella):

the volume of a cube is obtained by multiply the width * length * height. As it is a cube all of these sides are equal. So we can write x^3 = 2197 which means x * x * x = 2197. x is the cube root of 2197.

OpenStudy (welshfella):

( x is the length of the sides of the cube) A simple cube is 8 2*2*2 = 8 - therefore 2 is the cube root of 8.

OpenStudy (welshfella):

So we need to find a number which mutliplied by itself 3 times gives 2197. A scientific calculator will give you this. Have you got one?

OpenStudy (anonymous):

no sorry

OpenStudy (welshfella):

Are you on a PC?

OpenStudy (welshfella):

or laptop?

OpenStudy (welshfella):

A windows Pc will have a scientific calculator